OpenStudy (jjuden):

Which equation has intercepts at X(1, 0, 0), Y(0, -1, 0), and Z(0, 0, 2)?

OpenStudy (jjuden):

a x - y + 2z = 4 b 3x- 2y + 6z = 12 c 2x - 2y + z = 2 d 4x -4y + z = 4

ganeshie8 (ganeshie8):

plugin (1, 0, 0) into each of the options

OpenStudy (jjuden):

ok

ganeshie8 (ganeshie8):

which option(s) satisfy that point?

OpenStudy (jjuden):

still doing it hold on